DETECTION OF D_{2}H^{+} IN THE DENSE INTERSTELLAR MEDIUM

\underline{C. VASTEL}, T.G. PHILLIPS, California Institute of Technology, Downs Laboratory of Physics, MS 320-47, 1200 East California Boulevard, Pasadena, CA 91125, US; H. YOSHIDA, Caltech Submillimeter Observatory, 111 Nowelo Street, Hilo, HI 96720, US.

The 692 GHz para ground-state line of D_{2}H^{+} has been detected at the Caltech Submillimeter Observatory towards the pre-stellar core 16293E. The derived D_{2}H^{+} abundance is comparable to that of H_{2}D^{+}, as determined by observations of the 372 GHz line of ortho-H_{2}D^{+}. This is an observational verification of recent theoretical predictions (Roberts, Herbst & Millar 2003), developed to explain the large deuteration ratios observed in cold, dense regions of the interstellar medium associated with low mass pre-stellar cores and protostars. This detection confirms expectations that the multiply deuterated forms of H_{3}^{+} were missing factors of earlier models. The inclusion of D_{2}H^{+} and D_{3}^{+} in the models leads to predictions of higher values of the D/H ratio in the gas phase.